# Hybrid Reinforcement:When Reward Is Sparse, It’s Better to Be Dense

### v1

@@ Abstract
Post-training for reasoning of Large language models (LLMs) increasingly rely on verifiable rewards: deterministic checkers that provide 011 correctness signals. While reliable, such binary feedback is brittle—many tasks admit partially correct or alternative answers that verifiers under-credit, and the resulting all-or-nothing supervision limits learning. Reward models offer richer, continuous feedback, which can serve as a complementary supervisory signal to verifiers. We introduce HERO (Hybrid Ensemble Reward Optimization), a reinforcement learning framework that integrates verifier signals with reward-model scores in a structured way. HERO employs stratified normalization to bound reward-model scores within verifier-defined groups, preserving correctness while refining quality distinctions, and variance-aware weighting to emphasize challenging prompts where dense signals matter most. Across diverse mathematical reasoning benchmarks, HERO consistently outperforms RM-only and verifier-only baselines, with strong gains on both verifiable and hard-to-verify tasks. Our results show that hybrid reward design retains the stability of verifiers while leveraging the nuance of reward models to advance reasoning.
@@
Ping Yu at
Figure 1: Comparison of reward signals from different supervision sources. Blue dots indicate correct responses, while red crosses denote incorrect ones. The Reward Model (a) provides smooth but sometimes misaligned scores, occasionally assigning high values to incorrect responses and low values to correct ones. The Rule-based Method (b) enforces a strict 0-1 boundary: it rarely gives false positives, but due to its stringent criteria, many predictions that are actually correct receive a reward of 0 simply because they fail to pass the rule. HERO (c) uses the rule as a gate, which significantly reduces false positives. At the same time, by integrating the reward signal, HERO assigns higher reward scores to those cases that would have been false negatives under (b), resulting in more accurate and informative supervision.
@@ 1 Introduction
Reasoning lies at the heart of human intelligence, and increasingly, at the frontier of large language model (LLM) capabilities (Jaech et al., 2024; Guo et al., 2025; Zhang et al., 2025b). In tasks such as mathematical problems or generating proofs, reliable reasoning requires models to not only produce fluent text but also to generate logically consistent multi-step solutions that culminate in a verifiably correct outcome. Verifiable rewards implement this idea by running a deterministic checker—such as an exact numeric or string match, a unit test, or a symbolic equivalence check—on a candidate solution $y$ for input $x$. The checker either accepts or rejects the output, producing a sparse but unambiguous signal $r(x,y)\in\{0,1\}$, which reinforcement learning can propagate through the entire trajectory. Building on this principle, reinforcement learning from verifiable rewards (RLVR) (Chen et al., 2025b) uses these binary signals to train policies toward solutions that pass the checker. Recent systems—including OpenAI o1 and DeepSeek-R1—have advanced this paradigm at scale, leveraging verifier-grounded feedback to improve reasoning (Jaech et al., 2024; Guo et al., 2025; Zeng et al., 2025; Luo et al., 2025; Yang et al., 2024a).
@@
However, strict 011 verification is inherently coarse and brittle: many reasoning tasks allow for partially correct solutions, equivalent answers in alternative formats, or open-ended outputs that resist exact matching. In such cases, symbolic verifiers may under-credit valid solutions (false negatives) or fail to provide any useful signal (Ma et al., 2025; Huang et al., 2025a). Even when applicable, binary rewards induce sparsity: if all rollouts for a prompt receive the same label (all 0s or 1s), group-relative methods such as GRPO (Shao et al., 2024) yield zero relative advantage and thus no useful policy gradient, stalling policy improvement. This brittleness not only reduces sample efficiency but also skews optimization toward easier, strictly verifiable cases—leaving the hardest and most informative prompts underutilized. Our motivating analysis in Section 3.1 further highlights this limitation: on samples where answers are hard to verify, rule-based verifiers frequently fail to capture correctness. Reward-based models, in contrast, offer dense supervision by scoring responses on a continuum (Yang et al., 2024b; Liu et al., 2024; Zhang et al., 2025c; Lyu et al., 2025; Liu et al., 2025). Rather than collapsing all incorrect answers into the same category, they can capture nuanced quality differences such as partial correctness, clarity of reasoning steps, or proximity to the ground truth. This graded feedback enriches training signals, helping policies learn from partially correct reasoning paths and better allocate credit across diverse rollouts. However, naively combining these dense reward model signals with a binary verifier output often destabilizes training. Specifically, when the reward model’s continuous signals are naively blended with binary correctness checks, the resulting reward can become noisy or misaligned with the expected semantics of correctness. Figure 1 illustrates this tradeoff: while reward models offer smooth but misaligned signals, rule-based verifiers enforce correctness but lack nuance. Thus, it remains an open question of how to design an effective hybrid framework that preserves the reliability of verifiers while harnessing the richness of reward models?
@@
To address this challenge, we propose HERO (Hybrid Ensemble Reward Optimization), a reinforcement learning framework that integrates verifier-anchored and dense reward-model signals in a structured way. HERO integrates both to provide reliable yet informative supervision. HERO tackles the instability of naive blending through two key innovations. First, it introduces a stratified normalization scheme that bounds reward-model scores within verifier-defined correctness groups. This ensures that dense feedback refines learning only within the set of responses deemed correct by the verifier, preserving correctness guarantees while exploiting nuanced distinctions. Second, HERO employs a variance-aware weighting mechanism that adaptively adjusts the contribution of different prompts during training. Easy prompts, where most responses are uniformly correct or incorrect, contribute little additional learning signal and are down-weighted. In contrast, harder prompts—where candidate responses vary widely and reward-model scores provide valuable discrimination—are emphasized. These components allow HERO to overcome the brittleness of purely binary rewards and the noisiness of dense signals.
@@
We evaluate HERO on diverse math reasoning benchmarks that span three regimes: verifiable tasks where exact final-answer checking is possible, hard-to-verify tasks with partially correct or format-sensitive solutions, and mixed settings combining both. Across different LLM backbones, HERO consistently outperforms both RM-only and verifier-only baselines, in all three regimes. Notably, on hard-to-verify tasks, HERO achieves 66.3, which surpasses RM-only (54.6) by +11.7 points and verifier-only (57.1) by a dramatic +9.2 points. Ablations further confirm that anchoring dense signals to verifiable correctness and adaptively reweighting difficult prompts are both critical for stability and efficiency.
@@ 2 Preliminaries
Dense reward via reward modeling. Reward modeling learns a scalar function $r(x,y)$ that evaluates the quality of a response $y$ given a prompt $x$. Based on the Bradley–Terry model (Bradley & Terry, 1952), the reward function is trained on pairwise preference data by minimizing
@@
Group Relative Policy Optimization. GRPO (Shao et al., 2024) extends RLVR by optimizing over multiple responses per prompt rather than treating them independently. Instead of relying on a single trajectory, GRPO compares groups of candidate solutions and assigns relative advantages, which stabilizes learning and improves exploration. It also incorporates clipping (as in PPO) to prevent unstable updates and adds a KL penalty to keep the policy close to a reference model. This group-based formulation alleviates the gradient sparsity problem of pure verifier rewards and makes optimization more sample-efficient than standard PPO (Yu et al., 2025).
@@ 3 Methodology
Building on the preliminaries, we now examine how the two supervision paradigms – rule-based verifiers that provide sparse but precise correctness signals, and reward models that offer dense but potentially noisy preferences – behave on tasks where correctness is difficult to verify. Since the reliability of training hinges on the quality of supervision, understanding their respective strengths and weaknesses is crucial. To this end, we use the HardVerify_Math benchmark (Xu et al., 2025) as prompts, generate three responses per problem from Llama3.1-8B, Llama3.3-70B, and Qwen3-8B, and then evaluate verifier and reward model performance on these samples.
@@
To better understand the trade-offs among different verification approaches, we compare several representative verifiers. For rule-based verifiers, we consider math_reward.py from the verl library, math_verify module from verl, and the parse and verify functions from the Math-Verify library. In addition, we include more general verifiers that utilize a binary classifier trained to judge the correctness of answers, such as the TIGER-Lab/general-verifier (Ma et al., 2025).
@@
Table 1: Performance Comparison of Rule-based Verifier, LLM-as-Verifier, and Reward Model.
@@
LLM-as-verifier general-verifier 49.5 89.3 6.3 70.9
@@
We further examine how reward models behave on hard-to-verify samples. Since correctness is directly checkable, most reward models for mathematical reasoning are trained on verifiable samples (Yang et al., 2024b; Liu et al., 2024; Zhang et al., 2025c; Lyu et al., 2025; Liu et al., 2025). This raises the question: to what extent can such models generalize to tasks where correctness cannot be directly verified?
Here, we investigate this issue by analyzing the performance of a math-focused reward model (AceMath-7B-RM) on hard-to-verify tasks. We assess the model using different score thresholds generated by the reward model. As shown in Table 1, setting the threshold to 1 (i.e., considering predictions with RM $\geq$ 1 as correct) yields a high recall of 91.7% and broad overall coverage, substantially surpassing rule-based verifiers. However, this comes at the cost of lower precision. Increasing the threshold enhances precision but leads to a decrease in recall.
@@ 3.2 HERO: Hybrid Ensemble Reward Optimization
To address this, we propose stratified normalization, which explicitly bounds the continuous scores of RM within the symbolic structure imposed by the verifier. Formally, let $r_{\text{rule}}\in\{0,1\}$ denote the verifier output and $r_{\text{RM}}\in\mathbb{R}$ the reward-model score. We partition responses by $r_{\text{rule}}$ and apply group-wise min–max normalization to $r_{\text{RM}}$, yielding:
@@
Here $\alpha,\beta\in(0,1]$ control the allowable ranges for incorrect and correct groups, with $\epsilon>0$ preventing division by zero. Technically, we set this value relatively small so that the training dynamic is primarily led by rule-based rewards, and the reward from reward modeling is only supplementary. This design differs from traditional pure verifiable reward in the hard-to-verify samples and all-positive and all-negative groups, which do not provide the advantage over different rollouts.
This stratified normalization effectively embodies the hybrid approach: verifiers ensure the preservation of correctness semantics by constraining the score ranges, while reward models enhance the supervision by introducing gradations within each group. Incorrect responses are clearly distinguished from correct ones, and correct responses are prioritized based on their relative quality. In this manner, dense signals are anchored to symbolic correctness, mitigating the sparsity observed in pure RLVR.
@@
where $w_{\min}$ and $w_{\max}$ set the minimum and maximum weights, and $k$ controls the slope of the transition. In practice, we treat these as tunable hyperparameters; unless otherwise stated, we use $w_{\min}=0.5$, $w_{\max}=2.0$, and $k=5$, ensuring that difficult prompts are up-weighted by at most $2\times$ while trivial prompts retain at least half weight. The final shaped reward is
@@
This design operationalizes our intuition: ambiguous, high-variance prompts are emphasized because they reveal more about model weaknesses and reward-model sensitivity, while trivial, low-variance prompts are down-weighted to avoid wasting capacity. In doing so, the training process not only remains anchored to verifiable correctness through $\hat{r}$, but also allocates learning effort to the most challenging and informative parts of the data.
@@ 4.1 Experimental setup
A central question raised is whether reasoning skills acquired through RLVR on verifiable data can generalize to tasks whose correctness cannot be mechanically checked. To empirically examine this, we design three evaluation regimes: verifiable-only, hard-to-verify-only, and mixed. To evaluate learning under different types of supervision, we construct three training regimes based on subsets of the OpenMathReasoning (Moshkov et al., 2025) benchmark. For the verifiable-only regime, we sample 2,000 problems whose final answers can be deterministically validated using a rule-based math_verifier. For the hard-to-verify-only regime, we likewise sample 2,000 problems from OpenMathReasoning, which consists of the correct answer whose format is very complex (see Appendix A.2.2 for how we filter as well as some qualitative examples). These tasks supply dense preference signals and model-based verifier scores, but lack reliable binary labels from exact checking. Finally, in the mixed regime, we combine 1,000 verifiable and 1,000 hard-to-verify problems per epoch, enabling the policy to benefit simultaneously from robust exact-check supervision and nuanced feedback from unverifiable cases. Unless otherwise stated, mini-batches are stratified so that each epoch preserves the designated regime’s composition, and variants or prompts are randomly resampled to reduce overfitting to a single rendition.
@@
To evaluate the generalizability of our method across different backbone models, we conduct experiments using the following models of various model families and sizes: we use Qwen3-4B-Base (Yang et al., 2025) and Octothinker Hybrid 8B base mode (Wang et al., 2025). Motivated by stabilizing the RL training dynamic, we perform the cold start SFT on the base model (see Appendix A.2.1). All of the experiments of RL training start from the same SFT model.
@@
To provide a comprehensive comparison, we benchmark our hybrid-reward framework against both standard RL paradigms and stronger model-based references. As preliminary points of reference, we also report the performance of the base model and a cold-start SFT model, which serve to contextualize the impact of reinforcement learning itself. The main baselines are: (1) Reward model (RM)-only RL, which uses the AceMath-RM-7B reward model (Liu et al., 2024) to provide dense supervision; (2) Math verifier, which relies on binary, rule-based rewards, marking a sample as correct only if the normalized final answer matches the ground truth via math_verifier in the VERL repo—this emphasizes stability and reliability.
Evaluation for verifiable tasks.
We report pass@1 averaged over 8 seeds in Table 2. Following a standard decoding protocol, we use temperature $0.6$ and top-$p$ $0.95$, generate $N=8$ candidates per problem, and evaluate the first decoded output (pass@1). Reported numbers are means over seeds. Correctness is decided by math_verifier (normalized numeric/string match with task-specific post-processing). Benchmarks include MATH500 (Hendrycks et al., 2021), AMC (Li et al., 2024), Minerva (Lewkowycz et al., 2022), and Olympiad (He et al., 2024).
Evaluation for hard-to-verify tasks.
Since symbolic checkers cannot reliably provide binary labels for open-ended solutions, we adopt an LLM-as-a-judge protocol. Specifically, we use GPT-4o to compare model outputs against ground-truth answers. We evaluate using the HardVerify-Math benchmark (Xu et al., 2025), which consists of 250 samples. Based on the results in Section 3.1, we find that HardVerify-Math is not a particularly challenging filter, as using math_verify yields relatively good results. Therefore, to further evaluate performance on hard-to-verify reasoning tasks, we additionally collect the TextBookReasoning dataset (Fan et al., 2025) (see Appendix A.2.3 for more details).

### v3

@@ Abstract
Post-training for reasoning in large language models has increasingly relied on verifiable rewards: deterministic checkers that provide 011 correctness signals. While reliable, such binary feedback is brittle—many tasks admit partially correct or alternative answers that verifiers under-credit, and the resulting all-or-nothing supervision limits learning. Reward models offer richer, continuous feedback, which can serve as a complementary supervisory signal to verifiers. We introduce HERO (Hybrid Ensemble Reward Optimization), a reinforcement learning framework that integrates sparse verifier signals with dense reward model scores in a structured way. HERO employs stratified normalization to bound reward-model scores within verifier-defined groups, preserving correctness while refining quality distinctions, and variance-aware weighting to emphasize challenging prompts where dense signals matter most. Across diverse mathematical reasoning benchmarks, HERO consistently outperforms reward model-only and verifier-only baselines, with strong gains on both verifiable and hard-to-verify tasks. Our results show that hybrid reward design retains the stability of verifiers while leveraging the nuance of reward models to advance reasoning.
@@
Ping Yu at , Leitian Tao at
Figure 1: Comparison of reward signals from different supervision sources. Reward Models (a) provide smooth but sometimes misaligned scores, occasionally assigning high values to incorrect responses and low values to correct ones. Rule-based rewards (b) enforce a strict binary (0–1) boundary: they rarely give false positives, but due to their stringent criteria, many predictions that are actually correct receive a reward of 0 simply because they fail to pass the rule. HERO (c) uses the rule as a gate, which significantly reduces false positives. At the same time, by integrating the reward model signal, HERO assigns higher reward scores to those cases that would have been false negatives under (b), resulting in more accurate and informative supervision.
@@ 1 Introduction
Reasoning lies at the heart of human intelligence, and increasingly, at the frontier of large language model (LLM) capabilities (Jaech et al., 2024; Guo et al., 2025; Zhang et al., 2025b). In tasks such as mathematical problems or generating proofs, reliable reasoning requires models to generate logically consistent multi-step solutions that culminate in a verifiably correct outcome. Verifiable rewards implement this idea by running a deterministic checker—such as an exact numeric or string match, a unit test, or a symbolic equivalence check—on a candidate solution $y$ for input $x$. The checker either accepts or rejects the output, producing a sparse but unambiguous signal $r(x,y)\in\{0,1\}$, which reinforcement learning can propagate through the entire trajectory. Building on this principle, reinforcement learning from verifiable rewards (RLVR) (Chen et al., 2025b) uses these binary signals to train policies toward solutions that pass the checker. Recent systems—including DeepSeek-R1—have advanced this paradigm at scale, leveraging verifier-grounded feedback to improve reasoning (Guo et al., 2025; Zeng et al., 2025; Luo et al., 2025; Yang et al., 2024a).
@@
However, strict 011 verification is coarse and brittle: many reasoning tasks allow for partially correct solutions, equivalent answers in alternative formats, or open-ended outputs that resist exact matching. In such cases, symbolic verifiers may under-credit valid solutions (false negatives) or fail to provide any useful signal (Ma et al., 2025; Huang et al., 2025a). Even when applicable, binary rewards induce sparsity: if all rollouts for a prompt receive the same label (all 0s or 1s), group-relative methods such as GRPO (Shao et al., 2024) yield zero relative advantage and thus no useful policy gradient, stalling policy improvement. This brittleness not only reduces sample efficiency but also skews optimization toward easier, strictly verifiable cases–leaving the hardest and most informative prompts underutilized. Our motivating analysis in Section 3.1 further highlights this limitation: on samples where answers are hard to verify, rule-based verifiers frequently fail with correctness. Reward models, in contrast, offer dense supervision by scoring responses on a continuum (Yang et al., 2024b; Liu et al., 2024; Zhang et al., 2025c; Lyu et al., 2025; Liu et al., 2025). Rather than collapsing all incorrect answers into the same category, they capture nuanced quality differences such as partial correctness, clarity of reasoning steps, or proximity to the ground truth. This graded feedback enriches training signals, helping policies learn from partially correct reasoning paths and better allocate credit across diverse rollouts. However, naively combining these dense reward CT model signals with a binary verifier output often destabilizes training. Specifically, when the reward model’s continuous signals are naively blended with binary correctness checks, the resulting reward can become noisy or misaligned with the expected semantics of correctness. Figure 1 illustrates this tradeoff: while reward models offer smooth but misaligned signals, rule-based verifiers enforce correctness but lack nuance. Thus, it remains an open question how to design an effective hybrid framework that preserves the reliability of verifiers while harnessing the richness of reward models.
@@
To address this challenge, we propose HERO (Hybrid Ensemble Reward Optimization), a reinforcement learning framework that integrates verifier-anchored and dense reward-model signals to provide reliable yet informative supervision. HERO tackles the instability of naive blending through two key innovations. First, it introduces a stratified normalization scheme that bounds reward-model scores within verifier-defined correctness groups. This ensures that dense feedback refines learning only within the set of responses deemed correct by the verifier, preserving correctness guarantees while exploiting nuanced distinctions. Second, HERO employs a variance-aware weighting mechanism that adaptively adjusts the contribution of different prompts during training. Easy prompts, where most responses are uniformly correct or incorrect, contribute little additional learning signal and are down-weighted. In contrast, harder prompts—where candidate responses vary widely and reward-model scores provide valuable discrimination—are emphasized. These components allow HERO to overcome the brittleness of purely binary rewards and the noisiness of dense signals.
@@
We evaluate HERO on diverse math reasoning benchmarks that training with three regimes: training with easy-to-verify samples where exact final-answer checking is possible, hard-to-verify samples with partially correct or format-sensitive solutions, and mixed settings combining both. Across different LLM backbones, HERO consistently outperforms both RM-only and verifier-only baselines, in all three regimes. Notably, on hard-to-verify tasks evaluation based on Qwen-4B-Base, HERO achieves 66.3, which surpasses RM-only training (54.6) by +11.7 points and verifier-only training (57.1) by +9.2 points. Ablations further confirm that anchoring dense signals to verifiable correctness and adaptively reweighting difficult prompts are both critical for stability and efficiency.
@@ 2 Preliminaries
Dense reward via reward modeling. Reward modeling learns a scalar function $r(x,y)$ that evaluates the quality of a response $y$ given a prompt $x$. Based on the Bradley–Terry model (Bradley & Terry, 1952), the reward function is typically trained on pairwise preference data by minimizing
@@
Group Relative Policy Optimization. GRPO (Shao et al., 2024) extends RLVR by optimizing over multiple responses per prompt rather than treating them independently. Instead of relying on a single trajectory, GRPO compares groups of candidate solutions and assigns relative advantages, which stabilizes learning and improves exploration. It also incorporates clipping (as in PPO) to prevent unstable updates and adds a KL penalty to keep the policy close to a reference model. This group-based formulation helps alleviate the gradient sparsity problem of pure verifier rewards and makes optimization more sample-efficient than standard PPO (Yu et al., 2025).
@@ 3 Methodology
Building on the preliminaries, we now examine how the two supervision paradigms – rule-based verifiers that provide sparse but precise correctness signals, and reward models that offer dense but potentially noisy preferences – behave on tasks where correctness is difficult to verify. Since the reliability of training hinges on the quality of supervision, understanding their respective strengths and weaknesses is crucial. To this end, we use the HardVerify_Math benchmark (Xu et al., 2025), which focuses on challenging verification scenarios. This benchmark consists of 250 hard-to-verify math problems, including 115 manually selected Olympiad questions (He et al., 2024) and 10 MATH test set questions (Hendrycks et al., 2021) that are prone to false negatives due to complex answer formats, as well as 125 Big-Math questions (Albalak et al., 2025) with a Llama-3.1-8B (Dubey et al., 2024) pass rate below 0.05 and identified as difficult by human experts. To ensure a diverse range of response qualities, for each problem we generate answers using three different models: Llama-3.1-8B, Llama-3.3-70B, and Qwen3-8B. This results in a total of 750 responses, which we use to conduct the verifier analysis presented in Table 1.
@@
To better understand the trade-offs among different verification approaches, we compare several representative verifiers. For rule-based verifiers, we consider math_reward.py from the verl library, the math_verify module from verl, and the parse and verify functions from the Math-Verify library. In addition, we include a generative model-based verifier (TIGER-Lab/general-verifier (Ma et al., 2025)), which is specifically trained for chain-of-thought answer verification. This model has demonstrated strong performance and serves as an effective alternative to traditional rule-based methods.
@@
Table 1: Performance Comparison of Rule-based Verifier, LLM-as-Verifier, and Reward Models.
@@
Generative Model-based TIGER-Lab/general-verifier 49.5 89.3 6.3 70.9
@@
We further examine how reward models behave on hard-to-verify samples. Since correctness is directly checkable for verifiable data, most reward models for mathematical reasoning are trained on these verifiable samples (Yang et al., 2024b; Liu et al., 2024; Zhang et al., 2025c; Lyu et al., 2025; Liu et al., 2025). This raises the question: to what extent can such models generalize to tasks where correctness cannot be directly verified?
Here, we investigate this issue by analyzing the performance of a math-focused reward model (AceMath-7B-RM) on the same hard-to-verify tasks. We assess the model using different score thresholds given the scores generated. As shown in Table 1, setting the threshold to 1 (i.e., considering predictions with RM $\geq$ 1 as correct) yields a high recall of 91.7% and broad overall coverage, substantially surpassing rule-based verifiers. However, this comes at the cost of lower precision. Increasing the threshold enhances precision but leads to a decrease in recall.
@@ 3.2 HERO: Hybrid Ensemble Reward Optimization
To address this, we propose stratified normalization, which rescales the continuous scores of the reward model (RM) to align with the range used by traditional binary rule-based methods. Specifically, let $\{r_{\text{rule}}^{(i)}\}_{i=1}^{N}\subseteq\{0,1\}$ denote the rule-based verifier outputs and $\{r_{\text{RM}}^{(i)}\}_{i=1}^{N}\subseteq\mathbb{R}$ the corresponding reward-model scores for a group of $N$ rollouts. We first partition the responses according to $r_{\text{rule}}$, and then apply min–max normalization within each group to $r_{\text{RM}}$ resulting in::
@@
Here $\alpha,\beta\in(0,1]$ control the allowable ranges for incorrect and correct groups, with $\epsilon>0$ preventing division by zero. In practice, we set $\epsilon$ to relatively small value so that the training dynamics are primarily led by rule-based rewards, with the reward model’s contribution serving as a supplementary signal. Figure 1 (c) illustrates the effect of our hybrid reward design compared to $r_{\text{RM}}$ (a) and $r_{\text{rule}}$ (b) alone. This design notably differs from traditional pure verifiable rewards, especially for hard-to-verify samples and for groups where all responses are either positive or negative. In such cases, pure rule-based methods do not distinguish between different rollouts, providing no learning signal. As illustrated in Figure 1(c), HERO introduces reward differences within regions where the rule-based rewards are all 0 or all 1, thereby enabling meaningful gradient flow even when the rule-based verifier assigns the same outcome to all rollouts.
The stratified normalization in our hybrid approach is designed to provide the best of both worlds: verifiers ensure the preservation of correctness semantics by constraining the score ranges, while reward models enhance the supervision by introducing gradations within each group. Incorrect responses are clearly distinguished from correct ones, and correct responses are prioritized based on their relative quality. In this manner, dense signals are anchored to symbolic correctness, mitigating the sparsity observed in pure RLVR.
@@
where $w_{\min}$ and $w_{\max}$ set the minimum and maximum weights, and $k$ controls the slope of the transition. In practice, we treat these as tunable hyperparameters; unless otherwise stated, we use $w_{\min}=0.5$, $w_{\max}=2.0$, and $k=5$, ensuring that difficult prompts are up-weighted by at most $2\times$, while trivial prompts retain at least half weight. The final shaped reward is
@@
This design operationalizes our intuition: ambiguous, high-variance prompts are emphasized because they reveal more about model weaknesses and reward-model sensitivity, while trivial, low-variance prompts are downweighted to avoid wasting capacity. In doing so, the training process not only remains anchored to verifiable correctness through $\hat{r}$, but also allocates learning effort to the most challenging and informative parts of the data.
@@ 4.1 Experimental setup
A central question is whether reasoning skills acquired through RLVR on verifiable data can generalize to tasks whose correctness cannot be mechanically checked. To empirically examine this, we construct three distinct training datasets based on subsets of the OpenMathReasoning (Moshkov et al., 2025) benchmark: (1). easy-to-verify examples, (2). hard-to-verify examples, and (3). a mixture of the two. For easy-to-verify training data, we sample 2,000 problems whose final answers can be deterministically validated using a rule-based math_verifier (verl). For the hard-to-verify-only regime, we likewise sample 2,000 problems from OpenMathReasoning, where the correct answers have more flexible formats that make rule-based verification challenging (see Appendix A.2.2 for how we filter as well as some qualitative examples). For the mixed set, we combine 1,000 easy-to-verify and 1,000 hard-to-verify problems, allowing the policy to benefit from both robust exact-check supervision and nuanced feedback from unverifiable cases.
@@
To evaluate the generalizability of our method across different backbone models, we conduct experiments using the following models: we use Qwen3-4B-Base (Yang et al., 2025) and OctoThinker-8B-Hybrid-Base base (Wang et al., 2025). To stabilize RL training dynamics, we first perform supervised fine-tuning (SFT) on each base model as a cold start (see Appendix A.2.1 for details). All RL experiments are initialized from the same SFT checkpoint.
@@
As preliminary points of reference, we also report the performance of the base model and a cold-start SFT model. The main baselines are: (1) Reward model, which uses the AceMath-RM-7B reward model (Liu et al., 2024); (2) Rule-based verifier, which relies on binary, rule-based rewards, marking a sample as correct only if the normalized final answer matches the ground truth via math_verify (library) in the VERL repo (Sheng et al., 2025). Our method, HERO, is a hybrid approach that combines (1) and (2) into a single reward, making them the most appropriate baselines for comparison. We also compare HERO with a generative model-based verifier—TIGER-Lab/general-verifier—and with a large language model (Qwen2.5-7B-Instruct) directly prompted to act as a verifier, as detailed in the Appendix (see Table 8).
Evaluation.
Since our training data contains both easy-to-verify and hard-to-verify samples, we aim to evaluate whether the model can acquire generalizable reasoning abilities. To this end, we select six test sets: four in which all answers are easy to verify, and two in which the answers are hard to verify.
(1). Easy-to-verify testsets includes 4 benchmarks: MATH500 (Hendrycks et al., 2021), AMC (Li et al., 2024), Minerva (Lewkowycz et al., 2022), and Olympiad (He et al., 2024). We report pass@1 averaged over 8 seeds in Table 2. Following (Yang et al., 2024b), we use temperature $0.6$ and top-$p$ $0.95$, generate $N=8$ candidates per problem, and evaluate the first decoded output (pass@1). Reported numbers are means over seeds. Correctness is decided by math_verifier (normalized numeric/string match with task-specific post-processing). (2). Hard-to-verify testsets: We use temperature $0.6$ and top-$p$ $0.95$, generate $N=1$ candidate per problem. Since symbolic checkers cannot reliably provide binary labels for open-ended solutions, we adopt an LLM-as-a-judge protocol. Specifically, we use GPT-4o to compare model outputs against ground-truth answers. We evaluate using the HardVerify-Math benchmark (Xu et al., 2025), which consists of 250 samples. Based on the results in Section 3.1, we find that HardVerify-Math is not a particularly challenging filter, as using math_verify yields relatively good results. Therefore, to further evaluate performance on hard-to-verify reasoning tasks, we additionally collect the TextBookReasoning dataset (Fan et al., 2025) (see Appendix A.2.3 for more details).